OpenStudy (anonymous):

Beginning algebra: Mr. Crocker can pain his house by himself in 3 days. His son can paint the house by himself in 4 days. Find how long it takes to paint the house if they work together. The answer is 1 5/7 days. (mixed fraction answer)

OpenStudy (anonymous):

I would like to see the work. That's mainly what i'm looking for! Thanks

OpenStudy (anonymous):

Or a simple explanation...

OpenStudy (anonymous):

Never mind... I got it. im dumb

OpenStudy (anonymous):

\[\frac{1}{3}+\frac{1}{4}=\frac{1}{x} \]

OpenStudy (anonymous):

yup. that's exactly what I did. Thank!

OpenStudy (anonymous):

You are welcome.